### v1

# Optimizing a-DCF for Spoofing-Robust Speaker Verification

Abstract
Automatic speaker verification (ASV) systems are vulnerable to spoofing attacks such as text-to-speech. In this study, we propose a novel spoofing-robust ASV back-end classifier, optimized directly for the recently introduced, architecture-agnostic detection cost function (a-DCF). We combine a-DCF and binary cross-entropy (BCE) losses to optimize the network weights, combined by a novel, straightforward detection threshold optimization technique. Experiments on the ASVspoof2019 database demonstrate considerable improvement over the baseline optimized using BCE only (from minimum a-DCF of $0.1445$ to $0.1254$), representing $13\%$ relative improvement. These initial promising results demonstrate that it is possible to adjust an ASV system to find appropriate balance across the contradicting aims of user convenience and security against adversaries.
Index Terms:
a-DCF, spoofing-robust speaker verification
I Introduction
Automatic speaker verification (ASV) seeks to verify a claimed identity based on speech evidence [1]. Despite being a matured biometric technology, ASV is susceptible to spoofing through replay [2], text-to-speech [3] and other adversaries. To counter these risks, various countermeasures (CMs) or presentation attack detectors (PADs) [4], have been proposed for detecting spoofing. Typically CMs are developed in isolation from ASV, and the two subsystems are combined to enhance robustness of the biometric system against adversaries. The common strategies include tandem (cascade) combination [5] and fusion of detection scores [6] or embeddings [7, 8]. Besides systems constructed from combination of speaker and spoof detection subsystems, other studies consider deep end-to-end networks optimized to predict the decision directly from the speech inputs [9, 10].
Whether realized by combining two subsystems (ASV and CM), or by training a monolithic end-to-end model, a key consideration is how to evaluate and optimize the complete biometric authentication system. Any such system must trade-off between misses (false rejections) and false alarms (false acceptances)—proxies for user (in)convenience and security, respectively. The so-called tandem detection cost function (t-DCF) [5] can be used to evaluate the performance of cascaded combinations of ASV and CM under a Bayes-risk framework. Nonetheless, as t-DCF requires two sets of detection scores (and two detection thresholds) corresponding to the ASV and the CM subsystems, it is inapplicable for performance assessment of other types of architectures. To address these limitations, [11] proposed recently an architecture-agnostic detection cost function (a-DCF) applicable to a wider variety of system architectures. Unlike t-DCF, a-DCF requires a single detection score and a single threshold value only. This new metric is the primary metric in one of the two main tracks in the currently running ASVspoof 5 challenge [^1] https://www.asvspoof.org/ (referred June 28, 2024).
For the first time, we propose a spoofing-robust ASV back-end that is directly optimized for the new a-DCF metric (as shown in Fig. 1). As an evaluation metric based on hard error counting, the original a-DCF metric is non-differentiable and cannot be directly optimized using gradient descent schemes. Luckily, hard error counts can be “softened” (made differentiable). While direct optimization of the t-DCF was addressed in [12] through reinforcement learning, this approach is limited to the tandem architectures. By integrating a differentiable version of the a-DCF into model training, we aim to enhance the performance and robustness of speaker verification systems against spoofing attacks, ultimately contributing to the advancement of secure and reliable biometric authentication technologies. Besides integrating the a-DCF metric into model training, the threshold value is also jointly optimized during model training in the proposed method.
Figure 1: Proposed a-DCF optimization method for spoofing-robust speaker verification. The model $g(x)$ takes embeddings as input and outputs score distributions, from which error rates $\hat{P}_{\mathrm{miss}}^{\mathrm{tar}}$, $\hat{P}_{\mathrm{fa}}^{\mathrm{non}}$, and $\hat{P}_{\mathrm{fa}}^{\mathrm{spf}}$ are calculated. These error rates are then used to compute soft a-DCF, which serve as the loss function for model optimization through backpropagation, ensuring improved performance by minimizing the a-DCF.
II Detection Cost Functions
Conventional detection cost functions [13] are evaluation metrics used to assess the performance of binary classification systems, such as speaker verification and anti-spoofing. They focus on balancing the trade-off between false alarms and misses, assigning specific costs to each type of error. A binary classifier can be defined as a function $g:\mathcal{X}\to\rm I\!R$ that assigns a detection score $g(x)$ for a given input $x\in\mathcal{X}$. This score then can be converted into predicted binary class label $\hat{y}\in\{0,1\}$ by comparing it with a pre-set threshold $\tau$ by $\hat{y}=\mathrm{I}(g(x)>\tau)$. Here, $\mathrm{I}$ is the indicator function which returns $1$ if the condition is true (null hypothesis accepted) and 00 otherwise. In conventional ASV, $x$ consists of a pair of enrollment and test utterances (presented typically as a pair of speaker embeddings), and the null hypothesis is that the same speaker is present in both. In our spoofing-aware ASV models, $x$ additionally contains spoof embedding extracted from the test utterance by a CM model.
II-A DCF and a-DCF
Detection Cost Function (DCF) [13] is used for performance assessment of conventional ASV systems. It is defined as:
$\mathrm{DCF(\tau_{asv})}=C_{\mathrm{miss}}^{\mathrm{tar}}\cdot\pi_{\mathrm{tar}}\cdot P_{\mathrm{miss}}^{\mathrm{tar}}(\tau_{\mathrm{asv}})+C_{\mathrm{fa}}^{\mathrm{non}}\cdot\pi_{\mathrm{non}}\cdot P_{\mathrm{fa}}^{\mathrm{non}}(\tau_{\mathrm{asv}}),$ (1)
where $C_{\mathrm{miss}}^{\mathrm{tar}}$ is the cost of a miss, $C_{\mathrm{fa}}^{\mathrm{non}}$ is the cost of a false alarm, $P_{\mathrm{miss}}^{\mathrm{tar}}$ is the miss rate, $P_{\mathrm{fa}}^{\mathrm{non}}$ is false alarm rate, $\pi_{\mathrm{tar}}$ is the prior probability of the target and $\pi_{\mathrm{non}}=1-\pi_{\mathrm{tar}}$ is the prior probability of nontarget, and $\tau_{\mathrm{asv}}$ is the detection threshold. The DCF in Eq. (1) allows one to set the cost parameters according to the user’s application specification.
The architecture-agnostic detection cost function (a-DCF) [11] is a performance metric introduced for spoofing-aware speaker verification (SASV) task. It extends upon the DCF in Eq. (1) by including a third class (spoofing attack):
$\mathrm{a\!-\!DCF}(\tau_{\mathrm{sasv}})=$ $\;C_{\mathrm{miss}}^{\mathrm{tar}}\cdot\pi_{\mathrm{tar}}\cdot P_{\mathrm{miss}}^{\mathrm{tar}}(\tau_{\mathrm{sasv}})$ (2)
$+C_{\mathrm{fa}}^{\mathrm{non}}\cdot\pi_{\mathrm{non}}\cdot P_{\mathrm{fa}}^{\mathrm{non}}(\tau_{\mathrm{sasv}})$
$+C_{\mathrm{fa}}^{\mathrm{spf}}\cdot\pi_{\mathrm{spf}}\cdot P_{\mathrm{fa}}^{\mathrm{spf}}(\tau_{\mathrm{sasv}}),$
where $P_{\mathrm{fa}}^{\mathrm{spf}}$ is the false alarm for spoof trials, $\pi_{\mathrm{spf}}$ is the prior probability of spoofing, $C_{\mathrm{fa}}^{\mathrm{spf}}$ is the respective costs for spoof false alarms. $\tau_{\mathrm{sasv}}$ is the detection threshold for spoofing-aware speaker verification (SASV) system.
The miss rate ($P_{\mathrm{miss}}^{\mathrm{tar}}$) and the false alarm rates ($P_{\mathrm{fa}}^{\mathrm{non}}$, $P_{\mathrm{fa}}^{\mathrm{spf}}$) in the a-DCF metric are computed as follows:
$P_{\mathrm{miss}}^{\mathrm{tar}}(\tau_{\mathrm{sasv}})$ $=\frac{1}{N_{\mathrm{tar}}}\sum_{x\in\mathrm{tar}}\mathrm{I}(g(x)\leq\tau_{\mathrm{sasv}})$ (3)
$P_{\mathrm{fa}}^{\mathrm{non}}(\tau_{\mathrm{sasv}})$ $=\frac{1}{N_{\mathrm{non}}}\sum_{x\in\mathrm{non}}\mathrm{I}(g(x)>\tau_{\mathrm{sasv}})$
$P_{\mathrm{fa}}^{\mathrm{spf}}(\tau_{\text{sasv}})$ $=\frac{1}{N_{\mathrm{spf}}}\sum_{x\in\mathrm{spf}}\mathrm{I}(g(x)>\tau_{\mathrm{sasv}}),$
where $N_{\mathrm{tar}}$, $N_{\mathrm{non}}$, and $N_{\mathrm{spf}}$ represent the number of target, non-target, and spoof class trials, respectively.
II-B Differentiable a-DCF Metric
The indicator function used to compute false alarm and miss rates in Eq. (3) is non-differentiable. Therefore, the a-DCF metric can not be directly used as loss function to optimize an SASV system. In this study, we propose a differentiable (soft) version of a-DCF suitable for direct minimization using gradient descent. Similar approaches have been successfully applied to optimize DCF for text-dependent speaker verification [14] and t-DCF for the constrained class of tandem systems [15], but not the a-DCF. The a-DCF metric provides a more holistic evaluation by integrating the costs associated with speaker verification. This approach simplifies the benchmarking process for different ASV system architectures, as it eliminates the need for separate metrics for each component. By focusing on a single metric, the a-DCF metric promotes development of integrated systems that are optimized for spoofing-aware speaker verification.
The non-differentiable components in a-DCF metric are the miss rate ($P_{\text{miss}}$) and the two false alarm rates ($P_{\text{fa}}^{\text{non}}$, $P_{\text{fa}}^{\text{spf}}$). A common approach for ‘softening’ hard error counts [14, 15] is to replace the indicator function by a continuous, differentiable function. We adopt the following approximations:
$\hat{P}_{\mathrm{miss}}^{\mathrm{tar}}(\tau_{\mathrm{sasv}})$ $=\frac{1}{N_{\mathrm{tar}}}\sum_{x\in\mathrm{tar}}\sigma\left(\tau_{\mathrm{sasv}}-g(x)\right)$ (4)
$\hat{P}_{\mathrm{fa}}^{\mathrm{non}}(\tau_{\mathrm{sasv}})$ $=\frac{1}{N_{\mathrm{non}}}\sum_{x\in\mathrm{non}}\sigma\left(g(x)-\tau_{\mathrm{sasv}}\right)$
$\hat{P}_{\mathrm{fa}}^{\mathrm{spf}}(\tau_{\mathrm{sasv}})$ $=\frac{1}{N_{\mathrm{spf}}}\sum_{x\in\mathrm{spf}}\sigma\left(g(x)-\tau_{\mathrm{sasv}}\right),$
where $\sigma(z)=\frac{1}{1+e^{-z}}$ is the sigmoid function. Essentially, the error rates are approximated based on their distance from the threshold. The soft a-DCF loss is now defined as:
$\mathcal{L}_{\mathrm{a\!-\!DCF}}^{\mathrm{soft}}(\tau_{\mathrm{sasv}})=$ $\;C_{\mathrm{miss}}^{\mathrm{tar}}\cdot\pi_{\mathrm{tar}}\cdot\hat{P}_{\mathrm{miss}}^{\mathrm{tar}}(\tau_{\mathrm{sasv}})$ (5)
$+C_{\mathrm{fa}}^{\mathrm{non}}\cdot\pi_{\mathrm{non}}\cdot\hat{P}_{\mathrm{fa}}^{\mathrm{non}}(\tau_{\mathrm{sasv}})$
$+C_{\mathrm{fa}}^{\mathrm{spf}}\cdot\pi_{\mathrm{spf}}\cdot\hat{P}_{\mathrm{fa}}^{\mathrm{spf}}(\tau_{\mathrm{sasv}}).$
II-C Optimization of soft a-DCF
Algorithm 1 Optimization Algorithm
1:Input: Train data $\mathcal{D}_{\mathrm{trn}}$, Development data $\mathcal{D}_{\mathrm{dev}}$, Batch Size $B$, Number of Epochs $N$
2:Output: Best threshold $\hat{\tau}_{*}$, Model Parameters $\Theta_{*}$
3:Initialize: $\mathrm{min\_a\!-\!DCF}\leftarrow\infty$, $\tau\leftarrow 0.5$
4:for epoch $=1$ to $N$ do
5:    // optimize DNN weights
6:    for $X\leftarrow get\_minibatch(\mathcal{D}_{\mathrm{trn}},B)$ do
7:         $\mathcal{J}(\Theta_{i},\tau_{i})\leftarrow\left(\mathcal{L}_{\mathrm{a-DCF}}^{\mathrm{soft}}(X;\Theta_{i},\tau_{i})+\mathcal{L}_{\mathrm{BCE}}(X;\Theta_{i})\right)/2$
8:         $\Theta_{i+1}\leftarrow update\_network\_weights(\Theta_{i},\mathcal{J}(\Theta_{i},\tau_{i}))$
9:    end for
10:    $\Theta\leftarrow\Theta_{i+1}$
11:    // find optimal threshold using grid search
12:    $\tau_{i+1}\leftarrow\operatorname*{argmin}\limits_{\tau}\{\mathcal{L}_{\mathrm{a-DCF}}^{\mathrm{soft}}(\Theta,\tau)\}$ in $\mathcal{D}_{\mathrm{trn}}$
13:    // keep track of best model using dev data
14:    if $\mathcal{L}_{\mathrm{a-DCF}}^{\mathrm{soft}}(\tau_{i+1})<\mathrm{min\_a\!-\!DCF}$ in $\mathcal{D}_{\mathrm{dev}}$ then
15:         $\mathrm{min\_a\!-\!DCF}\leftarrow\mathcal{L}_{\mathrm{a-DCF}}^{\mathrm{soft}}(\tau_{i+1})$
16:         $\hat{\tau}\leftarrow\tau_{i+1}$
17:    end if
18:end for
19:return $\hat{\tau}_{*}$, $\Theta_{*}$
The algorithm 1 presents the training loop of a neural network towards its optimization, with both soft a-DCF ($\mathcal{L}_{\mathrm{a-DCF}}^{\mathrm{soft}}(X;\Theta_{i},\tau_{i})$ ) and BCE loss ( $\mathcal{L}_{\mathrm{BCE}}(X;\Theta_{i})$ ) functions. The BCE loss is the binary cross entropy loss defined as:
$\mathcal{L}_{\mathrm{BCE}}=-\frac{1}{N}\sum_{i=1}^{N}[y_{i}\log(\hat{y_{i}})+(1-y_{i})\log(1-\hat{y_{i}})]$ (6)
where $N$ represents the number of training samples. Essentially, the soft a-DCF loss is a loss function of application dependent a-DCF metric and useful for fine-tuning the system towards a desired miss-false alarm balance. The BCE loss, in turn, helps to separate target from the pool of the two negative classes. For these reasons, we use the combination of these two losses. The algorithm initializes a minimum a-DCF to be infinity and a fixed threshold $\tau$. Then, for each epoch, it optimizes the model weights by calculating the combined loss $\mathcal{J}(\Theta_{i},\tau_{i})$ from soft a-DCF and BCE losses and updating the weights of the neural network. This combined loss is defined as:
$J(\Theta_{i},\tau_{i})=\frac{1}{2}\bigg{(}\mathcal{L}_{\mathrm{a-DCF}}^{\mathrm{soft}}(\tau_{\mathrm{sasv}})+\mathcal{L}_{\mathrm{BCE}}\bigg{)}.$ (7)
The algorithm then conducts a grid search (Line 12 in Algorithm 1) for a value of $\tau$ that gives the optimal threshold according to the a-DCF loss on the training data. The threshold value serves as the independent variable in Eq. (5) of the a-DCF loss. Using the a-DCF loss (or its combination with BCE loss) as the objective function during model training involves minimizing its value, which depends on the threshold value. This relationship necessitates finding the optimal threshold value that minimizes the a-DCF loss at each epoch which is achieved through grid search. This algorithm tracks the best model parameters and threshold as scored on development data, thus updating the lowest a-DCF found so far and its corresponding parameters. At the end, it returns the best threshold $\hat{\tau}_{*}$ and model parameters $\Theta_{*}$.
III Experimental Setup, Results and Discussion
We perform experiments on the ASVspoof2019 logical access (LA)  [3] dataset. The description of the dataset and experimental details are discussed in the following subsections.
III-A Dataset
The ASVspoof2019 LA subset consists of three parts, training, development, and evaluation. Each partition has a disjoint set of speakers. The average duration of the utterances is $2$ to $3$ seconds. The spoofed utterances are generated using $19$ VC and TTS algorithms ($6$ types of attacks in the training and the development sets, $13$ in the evaluation set).
III-B Experimental Setup
To demonstrate our proposed optimization scheme, we consider the DNN-based embedding fusion strategy ’Baseline2’ described in [7], that was used as a baseline in the SASV2022 challenge. Thereafter, many other studies follow similar approach to perform SASV [16, 17]. We aim to show whether our proposed technique brings any performance improvement or not. This model concatenates three embeddings ($\textbf{e}=\left[\textbf{e}_{\text{enr}}^{\text{asv}},\textbf{e}_{\text{tst}}^{\text{asv}},\textbf{e}_{\text{tst}}^{\text{cm}}\right]$): two speaker (ASV) embeddings from the enrolment ($\textbf{e}_{\text{enr}}^{\text{asv}}$) and test ($\textbf{e}_{\text{tst}}^{\text{asv}}$) utterances extracted using ECAPA-TDNN [18], and one spoofing (CM) embedding ($\textbf{e}_{\text{tst}}^{\text{cm}}$) extracted from the test utterance using AASIST [19]. The concatenated embeddings are fed to a DNN model that consists of three fully connected hidden layers with leaky ReLU activation functions. The layers contain $256$, $128$, and $64$ neurons, respectively. Unlike baseline, which contains two neurons with softmax, our model’s final layer contains a single neuron with a sigmoid activation function. This is because of our a-DCF optimization requires only one output score.
We consider four back-end models that are otherwise identical but optimized differently. These methods are:
  • S1: corresponds to the Baseline2 system provided by SASV challenge organizers (optimized with CE);
  • S2: optimized using soft a-DCF loss with, $\tau_{\mathrm{sasv}}=0.5$;
  • S3: optimized using soft a-DCF loss + BCE, with $\tau_{\mathrm{sasv}}=0.5$;
  • S4: optimized using soft a-DCF loss + BCE, including threshold optimization, as detailed in Algorithm 1.
The cost values $C_{\mathrm{miss}}^{\mathrm{tar}}$, $C_{\mathrm{fa}}^{\mathrm{non}}$, and $C_{\mathrm{fa}}^{\mathrm{spf}}$ in the soft a-DCF calculation are set as $1$, $10$, and $20$, respectively, while the prior values $\pi_{\mathrm{tar}}$, $\pi_{\mathrm{non}}$, and $\pi_{\mathrm{spf}}$ are set as $0.9$, $0.05$, and $0.05$. The higher relative cost value for spoof false alarms ($C_{\mathrm{fa}}^{\mathrm{spf}}=20$) is chosen based on the importance of spoofing robustness in security-sensitive domains, such as banking, where preventing unauthorized access is crucial. This reflects the emphasis to penalize false acceptance of spoof attempts to ensure the integrity and security of such systems. Additionally, the high target prior ($\pi_{\mathrm{tar}}=0.9$) indicates that the system is designed under the assumption that most interactions are legitimate. This is practical for applications where legitimate access attempts are frequent, and it minimizes the disruption to bonafide users by reducing the likelihood of false rejections. Except training, we considered the a-DCF metric given in Eq. 2 in all the experiments for performance evaluation.
TABLE I: Baseline system min a-DCF values in the development set
Batch size 24 48 96 192 384 768 1024
min a-DCF 0.1437 0.1308 0.1268 0.1346 0.1297 0.1251 0.1234
III-C Baseline (S1)
We firstly trained a baseline (S1) with cross entropy loss for various batch sizes, resulting to minimum a-DCF values shown in Table I. The results indicate that the best performance was achieved with a batch size of $1024$. We fix this value for the remainder of the experiments.
TABLE II: Comparison of S1 and S2 in development set. Batch size is selected as 1024.
System min a-DCF
S2 0.1355
S3 0.1182
III-D Proposed a-DCF metric optimization (S2, S3 and S4)
We next consider models S2 and S3, beginning with the hard-coded detection threshold of $\tau_{\mathrm{sasv}}=0.5$ in Eq. (5). Model selection is based on the lowest a-DCF value at a threshold of $0.5$ in the development set. The soft a-DCF only system (S2) attains minimum a-DCF of $0.1355$ and the combination of the soft a-DCF and BCE (S3) achieves minimum a-DCF of $0.1182$ as summarized in Table II.
TABLE III: Comparison of S1, S2, S3, and S4 in development and evaluation set. The results are given using 1024 as batch size.
Baseline (S1) S2 S3 S4
Dev 0.1234 0.1355 0.1182 0.1109
Eval 0.1445 0.2352 0.1398 0.1254
The results of S2 and S3 were based on hard-coded arbitrary detection threshold of $\tau_{\mathrm{sasv}}=0.5$. We now also optimize $\tau_{\mathrm{sasv}}$ using the complete procedure (Algorithm 1). Joint optimization of the DNN parameters and the threshold leads to noticeable improvements, as expected. Specifically, S4 achieved a minimum a-DCF of $0.1109$ on the development set and $0.1254$ on the evaluation set, outperforming S1, S2, and S3 as Table III indicates.
Figure 2: The plot illustrates the a-DCF values with respect to the threshold for the four different systems. The curves represent the a-DCF values for Baseline (S1), S2, S3, and S4 systems across various threshold values. Circular markers indicate the a-DCF values at a fixed threshold of $0.5$ for each system. Cross markers denote the minimum a-DCF values and their corresponding threshold points. The star marker represents the optimized threshold and the corresponding a-DCF value proposed for the S4 system.
Fig. 2 illustrates the a-DCF curves obtained for the four different systems (S1, S2, S3, and S4) as a function of the threshold. For systems S2 and S3, the a-DCF values at both the fixed threshold value of $0.5$ and their respective minimum points are depicted on the graph. For the S4 system, that includes threshold optimization, the a-DCF value corresponding to the optimized threshold is marked with a star. As can be seen from the graph (and supported by the results), the S4 model provides the lowest a-DCF value.
III-E Impact of a-DCF Parameter
Figure 3: Target vs Nontarget and Target vs Spoof DET curves for three different a-DCF parameter settings. The SV-EER values for Setting1 (represented by the solid blue line), Setting1 (soft a-DCF + BCE) (represented by the dotted blue line), Setting2, and Setting3 are 13.97%, 8.44%, 7.52%, and 49.13%, respectively.
Our final analysis addresses the impact of the a-DCF parameters. To this end, we introduced three different set of a-DCF parameters as:
  • Setting1 (soft a-DCF) contains only a-DCF loss. In this case, the cost values $C_{\mathrm{miss}}^{\mathrm{tar}}$, $C_{\mathrm{fa}}^{\mathrm{non}}$, and $C_{\mathrm{fa}}^{\mathrm{spf}}$ are set to $1$, $10$, and $20$, respectively. The prior values $\pi_{\mathrm{tar}}$, $\pi_{\mathrm{non}}$, and $\pi_{\mathrm{spf}}$ are set to $0.9$, $0.05$, and $0.05$.
  • Setting1 (soft a-DCF + BCE) contains both a-DCF and BCE losses. The cost and the prior values are the same as in the previous case.
  • Setting2 (soft a-DCF) contains only a-DCF loss. In this case, the cost values $C_{\mathrm{miss}}^{\mathrm{tar}}$, $C_{\mathrm{fa}}^{\mathrm{non}}$, and $C_{\mathrm{fa}}^{\mathrm{spf}}$ are all set to $1$. The prior values $\pi_{\mathrm{tar}}$, $\pi_{\mathrm{non}}$, and $\pi_{\mathrm{spf}}$ are set to $0.5$, $0.5$, and 00, respectively.
  • Setting3 (soft a-DCF) contains only a-DCF loss. The same cost values are used as in the previous setting, while $0.5$, 00 and $0.5$ priors are used for $\pi_{\mathrm{tar}}$, $\pi_{\mathrm{non}}$, and $\pi_{\mathrm{spf}}$, respectively.
The threshold optimization method was included for all these cases. Fig. 3 shows target vs. nontarget and target vs. spoof DET profiles for three different sets of a-DCF parameters. The figure clearly indicates that the choice of a-DCF parameters significantly impacts both DET profiles. Comparing setting1’s soft a-DCF alone with the soft a-DCF combined with BCE, the former generally has a lower miss rate. For setting2 and setting3, changes in the priors cause notable differences in the DET curves. For example, setting $\pi_{\mathrm{spf}}$ to 00 (setting2), assuming no spoof non-targets, shows the best performance for target vs. nontarget trials across all operating points. However, when the non-target prior is set to 00 (setting3), the performance for target vs. nontarget trials is found to be inferior, as expected. These trends are reversed for target vs. spoof trials.
IV Conclusion
In this study, we demonstrated that a spoofing-aware speaker verification system can be effectively optimized for the new a-DCF metric. By incorporating both the a-DCF and binary cross-entropy losses in the optimization process, we fine-tuned the network weights, which led to improved performance over the baseline system. Importantly, the usage of a novel threshold optimization technique further enhanced the system’s capabilities, resulting in a significant improvement in a-DCF values. These findings highlight the potential for adjusting ASV systems to achieve an optimal balance between user convenience and security against to evolving spoofing attacks.